# ROCKET-2: Steering Visuomotor Policy via Cross-View Goal Alignment

### v1

@@ Abstract
We aim to develop a goal specification method that is semantically clear, spatially sensitive, and intuitive for human users to guide agent interactions in embodied environments. Specifically, we propose a novel cross-view goal alignment framework that allows users to specify target objects using segmentation masks from their own camera views rather than the agent’s observations. We highlight that behavior cloning alone fails to align the agent’s behavior with human intent when the human and agent camera views differ significantly. To address this, we introduce two auxiliary objectives: cross-view consistency loss and target visibility loss, which explicitly enhance the agent’s spatial reasoning ability. According to this, we develop ROCKET-2, a state-of-the-art agent trained in Minecraft, achieving an improvement in the efficiency of inference 3×3$\times$3 × to 6×6$\times$6 ×. We show ROCKET-2 can directly interpret goals from human camera views for the first time, paving the way for better human-agent interaction. The project page is available at https://craftjarvis.github.io/ROCKET-2/.
[…]
Yitao Liang
Shaofei Cai, Zhancun Mu <{caishaofei, muzhancun}@stu.pku.edu.cn>, Anji Liu <liuanji@cs.ucla.edu>, Yitao Liang <yitaol@pku.edu.cn>
Figure 1: Cross-view Goal Alignment as a Human-Friendly Goal Specification Method. The target object can be specified using a segmentation mask from the human’s camera view, while the agent learns to align with human intent and take actions based on its own observations. Visual landmarks serve as key cues to establish spatial relationships across camera views.
@@ 1 Introduction
Learning an agent to achieve desired goals is a long-standing challenge in the field of embodied intelligence, with significant implications for the development of robots (Brohan et al., 2022, 2023; Jang et al., 2022) and virtual players (Wang et al., 2023b, c, a). A key challenge is to find goal representations that are (i) flexible for human users to specify and (ii) expressive and precise to capture as many tasks as possible. Most current approaches address only one of these aspects. For example, traditional works (Brohan et al., 2022; Driess et al., 2023; Lynch et al., 2023) focus on training agents to follow language instructions. As pointed out in Sundaresan et al. (2024); Cai et al. (2024b), while language is intuitive, it relies on numerous prepositions to express spatial relationships, which can be vague and inefficient. Furthermore, language also suffers from the generalization problem of novel visual concepts (Cai et al., 2023). Realizing these limitations, some works attempted to introduce visual modalities into goal representations. For example, Sundaresan et al. (2024) employs hand-drawn target layouts in robot manipulation environments to represent human intent; Gu et al. (2023) uses end-effector trajectory sketches for fine-grained control of robot arms; and ROCKET-1 (Cai et al., 2024b) specifies the objects to interact with by applying segmentation masks to the agent’s observations. These methods have greatly improved the expressiveness of spatial relationships and generalization across tasks. However, both trajectory sketches and object segmentation are closely tied to the agent’s current observation, causing issues in partially observable 3D worlds. These include: (i) goals need to be generated in real-time as the agent’s camera view changes; (ii) goals cannot be specified when the target is occluded.
@@
To strike a balance between expressiveness and flexibility, we propose an innovative and user-friendly cross-view goal specification method. It allows human users to specify the target object using segmentation masks from their own camera view, rather than the agent’s camera view. The agent is then trained to align with human intent and take actions based on its own observations via imitation learning. Decoupling the goal specification from the camera view of the agent will significantly enhance the efficiency of human-agent interaction. However, the partial observability of open worlds makes aligning goals across camera views challenging. This involves handling occlusion, geometric deformation, and the distinction of objects of similar look. In Figure 1, we show an agent in the left corner and a human player standing on a farmland. The human intends to command the agent to hunt a sheep near the house, even though the agent cannot initially observe the target sheep. To achieve this, the agent must establish spatial relationships using shared visual landmarks between the human’s camera view and its own. We find that relying solely on a behavior cloning loss is insufficient.
@@
To address these challenges, we highlight an important property of behavior datasets (Cai et al., 2024b): The target object remains consistent across camera views in an interaction trajectory. Motivated by this, we propose two auxiliary objective functions: cross-view consistency loss and target visibility loss, to explicitly enhance the agent’s ability to align goals across camera views. Specifically, cross-view consistency loss requires the agent to accurately predict the target object’s centroid point w.r.t. its camera view, while target visibility loss helps the agent determine whether the target object is occluded. To further leverage temporal consistency, we use a causal Transformer (Vaswani et al., 2017) in the architecture to model the relationship between past predictions and the current observation. It encourages the agent to maintain tracking even when the target is occluded. By combining these auxiliary losses with behavior cloning loss, we develop ROCKET-2, a state-of-the-art agent in Minecraft. Our experiments show that ROCKET-2 can autonomously track the target object as the camera changes, eliminating the need for SAM’s (Ravi et al., 2024) real-time semantic segmentation, speeding up inference 3×3$\times$3 × to 6×6$\times$6 × compared to ROCKET-1. To our knowledge, we are the first to demonstrate that agents can interpret intentions from a human’s camera view and make decisions to achieve expected goals in the embodied world. Extensive visualization experiments and case studies offer deeper insights into its behavior.
@@
Our contributions are threefold: (1) We introduce a user-friendly interface that allows humans to specify goals using segmentation masks from their camera view, paving the way for better human-agent interaction. (2) We propose cross-view consistency loss and target visibility loss to explicitly enhance the agent’s ability to align targets across camera views and improve its steerability. (3) We train ROCKET-2, an agent that autonomously tracks targets, eliminating the need for real-time goal segmentation and significantly speeding up inference.
@@ 2 Related Works
We address policy learning in partially observable 3D worlds (Savva et al., 2019; Cai et al., 2024a), where the agent receives only egocentric images rather than the full environmental state. Since the movement of the camera view is part of the action space, the policy must actively explore to locate key objects, a common challenge in visual navigation (Savva et al., 2019) and FPS games (Pearce and Zhu, 2022). While some methods (Huang et al., 2023; Jiang et al., 2024) incorporate 3D point clouds to provide global context, such data is often unavailable. More commonly, memory-based architectures integrate historical observations to build implicit 3D perception. For example, Zhao et al. (2023); Gadre et al. (2022) use RNN to avoid redundant exploration, and Baker et al. (2022) employs TransformerXL to retain achievements and complete the long-horizon diamond challenge in Minecraft (Guss et al., 2019). Building on this, we require the policy to align objects across camera views to ensure consistent tracking of targets as the camera view changes. Although cross-view alignment has been explored in computer vision for tasks like BEV segmentation (Borse et al., 2023) and pedestrian re-identification (Xu et al., 2019), we are the first to study its application in open-world decision-making.
@@
GCIL refers to algorithms that optimize conditional policies through imitation learning, primarily using behavior cloning loss (Pomerleau, 1988). The policy conditions can take various forms, such as language (Brohan et al., 2022, 2023; Lynch et al., 2023), images (Majumdar et al., 2022; Lifshitz et al., 2023; Sundaresan et al., 2024), videos (Cai et al., 2023, 2024c), or trajectory sketches (Wang et al., ; Gu et al., 2023). Compared to traditional imitation learning, GCIL provides a more explicit target during training, reducing the complexity of modeling the entire behavior space and making the policy steerable during inference.
Humans often use language to express their intentions, leading many studies (Brohan et al., 2022; Padalkar et al., 2023; Driess et al., 2023; Wang et al., 2023c) to focus on learning language-conditioned policies. However, language goals are often ambiguous and struggle to capture spatial details (Cai et al., 2024b; Gu et al., 2023). In tasks like navigation (Majumdar et al., 2022) and object manipulation (Wang et al., ), image-based goal modalities have been explored, where users provide target images to guide the policy. While effective at conveying spatial information, images often over-specify details, making policies sensitive to irrelevant factors like lighting, object appearance, or background textures. Sundaresan et al. (2024) addresses these issues by replacing the image with hand-drawn sketches. However, generating sketches that align with the current state and the desired goal is non-trivial. Gu et al. (2023) further proposes using trajectory sketches for finer control and better generalization across tasks, but this approach is not applicable in partially observable 3D worlds.
ROCKET-1 (Cai et al., 2024b) tackles interaction problems in 3D worlds by training a visuomotor policy to identify interaction targets based on semantic segmentations in the visual context. While it resolves traditional goal images’ ambiguity and generation challenges, it relies on SAM-2 (Ravi et al., 2024) to track goals and segment them during inference, severely limiting real-time performance. We propose a cross-view segmentation-conditioned policy that, unlike ROCKET-1, enables the policy to align the goal across camera views by itself and removes the need for real-time segmentation.
Hindsight Trajectory Relabeling
There are two main approaches to collecting labeled trajectory data: (1) providing instructions for contractors to collect trajectories in real-time, ensuring a causal link between actions and labels but incurring high costs and limited scalability; and (2) gathering large amounts of trajectories and generating labels through post-processing, known as hindsight trajectory relabeling. While the first approach (Lynch et al., 2023; Padalkar et al., 2023) produces higher-quality data, its cost constraints have led most research to adopt the second, more scalable one. Andrychowicz et al. (2017) was the first to reinterpret a trajectory’s behavior using its final frame, greatly improving data utilization and inspiring subsequent research on goal-conditioned policies. Lifshitz et al. (2023) extended the approach by using the last 16 frames of a trajectory as the more expressive goal. Sundaresan et al. (2024) introduced hand-drawn sketches as a goal modality, reducing semantic ambiguity. Gu et al. (2023) converted the robotic end-effector moving sketch to a 2D image as the goal, providing richer procedural details. Cai et al. (2024b) introduced backward trajectory relabeling, which first identifies interaction objects and events, then utilizes object tracking models (Ravi et al., 2024) to generate frame-level segmentations. This data supports training segmentation-conditioned policies. Our paper explores using this dataset to train policies for cross-view goal alignment.
3 Methods
Figure 2: ROCKET-2 Architecture. It consists of three parts: (1) a non-causal transformer for spatial fusion, which establishes the relationship between the agent’s and human’s camera views; (2) a causal transformer for temporal fusion, ensuring consistency for goal tracking; (3) a decoder module, made of a feedforward neural network (FFN), which predicts goal-related visuals cues and actions.
@@ 3 Methods
Our goal is to learn a goal-conditioned visuomotor policy, which allows humans to specify goal objects for interaction using semantic segmentation across camera views. Formally, we aim to learn a policy $\pi_{\text{cross}}(a_{t}|o_{1:t},\{o_{g},m_{g}\},c_{g})$, where $a_{t}$ represents the action at time $t$, $c_{g}$ denotes the type of interaction. In the Minecraft environment, an action corresponds to raw mouse and keyboard inputs. $o_{t}\in\mathbb{R}^{H\times W\times 3}$ denotes the environment observation at time $t$, and $o_{g}\in\mathbb{R}^{H\times W\times 3}$ represents an observation of the local environment from a specific camera view. Generally, $o_{g}$ and $o_{t}$ have some visual content overlap. $m_{g}\in\{0,1\}^{H\times W\times 1}$ is a segmentation mask for $o_{g}$, highlighting the target object within the camera view $o_{g}$. During inference, users select a view $o_{g}$ containing the desired object from historical observations returned by the environment and generates its corresponding semantic segmentation $m_{g}$. To train such visuomotor policy, we assume access to a dataset $\mathcal{D}_{\text{cross}}=\{c^{n},(o_{t}^{n},a_{t}^{n},m_{t}^{n})_{t=1}^{L(n)}\}_{n=1}^{N}$ consisting of $N$ successful demonstration episodes, $L(n)$ is the length of episode $n$. Within each episode, if $m_{t}$ is non-empty, all $(o_{t},m_{t})$ pairs indicate the same object. Consequently, we can arbitrarily pick one observation frame as the goal view condition for the entire trajectory.
@@
Without loss of generality, we use the Minecraft world as an example to illustrate the data generation process. Manually collecting datasets that meet the requirements is highly expensive. Thus, we employ the backward trajectory relabeling technique proposed in Cai et al. (2024b) to automate the annotation of the OpenAI Contractor Dataset (Baker et al., 2022), which consists of free-play trajectories from human players: $\mathcal{D}_{\text{raw}}=\{(o_{t}^{n},a_{t}^{n})_{t=1}^{L(n)}\}_{n=1}^{N}$. Specifically, for any given episode $n$, we first detect all frames $o^{n}_{j}$ where interaction events occur, identify the interaction type $c^{n}_{j}$, and localize the interacted object near frame $j$ using bounding boxes and point-based prompts. The SAM-2 (Ravi et al., 2024) model is then employed to generate the segmentation mask $m^{n}_{j}$ for the object. Starting from frame $j$, we traverse the trajectory backward and use the SAM-2 model to continuously generate segmentation masks for the object in real-time until either a new interaction event is encountered or a maximum tracking length is reached. Let $i$ denote the end frame. The resulting trajectory clip is then added to the training dataset: $\mathcal{D}_{\text{cross}}\leftarrow\mathcal{D}_{\text{cross}}\cup\{c_{j},(o^{n}_{t},a^{n}_{t},m^{n}_{t})_{t=i}^{j}\}.$ This ensures that every extracted clip is associated with a consistent interaction intent. The generated data encompasses the fundamental interaction types in Minecraft, including use, break, approach, craft, and kill entity. Among these, approach is a unique event, identified by detecting trajectory clips where the displacement exceeds a specified threshold. The object located at the center of the clip’s final frame is designated as the goal of the approach event.
@@
It is worth noting that incorporating the historical observation sequence $o_{1:t-1}$ as input is essential, especially when there is limited shared visual content between $o_{t}$ and $o_{g}$. This historical sequence acts as a smooth bridge to facilitate alignment. Since the goal object represented by the segmentation corresponds to the target of the policy’s interaction, this auxiliary task aligns the policy’s actions with its visual focus, effectively improving task performance.
@@
Due to the partial observability in 3D environments, it is common for target objects in interaction trajectories to disappear from the field of view and reappear later. During such intervals, the segmentation mask for the missing object is empty. To leverage this information, we propose training the model to predict whether the target object is currently visible, formulated as: $\pi_{\text{cross}}(v_{t}\mid o_{1:t},\{o_{g},m_{g}\})$, where $v_{t}$ is a binary indicator for empty segmentation masks. On the one hand, accurately predicting object visibility helps the policy better match the target object, avoiding a simple appearance similarity measurement between two frames. On the other hand, visibility information guides the policy to make more reasonable decisions, such as confidently approaching the goal object when it is visible or actively adjusting its camera to explore when the target is absent.
@@
Lets a training trajectory $n$ be denoted as $(c_{g},\{o_{t},m_{t}\}_{t=1}^{L(n)})$. A cross view index $g$ is sampled from $\{i|i\in[1,L(n)],m_{i}\neq\phi\}$. We resize all visual observations and their segmentation masks to $224\times 224$. For encoding the visual observation $o_{t}$, we utilize a DINO-pretrained (Caron et al., 2021) 3-channel ViT-B/16 (Dosovitskiy et al., 2020) (16 is the patch size), which outputs a token sequence of length 196, denoted as $\{\hat{o}_{t}^{i}\}_{i=1}^{196}$. Similarly, we encode the segmentation mask $m_{t}$ using a 1-input-channel ViT-tiny/16, yielding $\{\hat{m}_{t}^{i}\}_{i=1}^{196}$. The ViT-base/16 encoder is frozen during training for efficiency, while the ViT-tiny/16 is trainable. To ensure spatial alignment, we fuse the cross-view condition $(o_{g},m_{g})$ by concatenating the feature channels:
@@
Given the ability of self-attention mechanisms to capture spatial details across views, we concatenate the token sequences from two views into a single sequence of length 392. A non-causal Transformer encoder module is applied (Vaswani et al., 2017) for spatial fusion, obtaining a frame-level representation $x_{t}$:
@@
Subsequently, we leverage a causal TransformerXL (Dai et al., 2019) architecture to capture temporal information across the sequence:
@@
Finally, a simple linear layer maps $f_{t}$ to predict the action $\hat{a}_{t}$, centroid $\hat{p}_{t}$, and visibility $\hat{v}_{t}$. The loss function for episode $n$ is defined as:

### v2

@@ Abstract
We aim to develop a goal specification method that is semantically clear, spatially sensitive, domain-agnostic, and intuitive for human users to guide agent interactions in 3D environments. Specifically, we propose a novel cross-view goal alignment framework that allows users to specify target objects using segmentation masks from their camera views rather than the agent’s observations. We highlight that behavior cloning alone fails to align the agent’s behavior with human intent when the human and agent camera views differ significantly. To address this, we introduce two auxiliary objectives: cross-view consistency loss and target visibility loss, which explicitly enhance the agent’s spatial reasoning ability. According to this, we develop ROCKET-2, a state-of-the-art agent trained in Minecraft, achieving an improvement in the efficiency of inference 3×3$\times$3 × to 6×6$\times$6 × compared to ROCKET-1. We show that ROCKET-2 can directly interpret goals from human camera views, enabling better human-agent interaction. Remarkably, ROCKET-2 demonstrates zero-shot generalization capabilities: despite being trained exclusively on the Minecraft dataset, it can adapt and generalize to other 3D environments like Doom, DMLab, and Unreal through a simple action space mapping. The project page is available at https://craftjarvis.github.io/ROCKET-2/.
Figure 1: Powered by cross-view goal specification, we are the first to show that AI agents can complete complex tasks such as building a bridge and damaging the dragon in Minecraft. In addition, it demonstrates an impressive zero-shot generalization to other 3D games.
@@ 1 Introduction
Learning an agent to achieve desired goals is a long-standing challenge in the field of embodied intelligence, with significant implications for the development of robots (Brohan et al., 2022, 2023; Jang et al., 2022) and virtual players (Wang et al., 2023c, d, b). A key challenge is to find goal representations that are (i) flexible for human users to specify and (ii) expressive and precise to capture as many tasks as possible. Most current approaches address only one of these aspects. For example, traditional works (Brohan et al., 2022; Driess et al., 2023; Lynch et al., 2023) focus on training agents to follow language instructions. As pointed out in Sundaresan et al. ; Cai et al. (2025a), while language is intuitive, it relies on numerous prepositions to express spatial relationships, which can be vague and inefficient. Furthermore, language also suffers from the generalization problem of novel visual concepts (Cai et al., 2023). Realizing these limitations, some works attempted to introduce visual modalities into goal representations. For example, Sundaresan et al. employs hand-drawn target layouts in robot manipulation environments to represent human intent; Gu et al. (2023) uses end-effector trajectory sketches for fine-grained control of robot arms; and ROCKET-1 (Cai et al., 2025a) specifies the objects to interact with by applying segmentation masks to the agent’s perception. These methods greatly improved the expressiveness of spatial relationships and generalization across tasks. However, both trajectory sketches and object segmentation are closely tied to the agent’s current observation, causing issues in partially observable 3D worlds. These include: (i) goals need to be generated in real-time as the agent’s camera view changes; (ii) goals cannot be specified when the target is occluded.
@@
To strike a balance between expressiveness and flexibility, we propose an innovative and user-friendly cross-view goal specification method. It allows human users to specify the target object using segmentation masks from their own camera view, rather than the agent’s camera view. The agent is then trained to align with human intent and take actions based on its own observations via imitation learning. Decoupling the goal specification from the camera view of the agent will significantly enhance the efficiency of human-agent interaction. However, the partial observability of open worlds makes aligning goals across camera views challenging. This involves handling occlusion, geometric deformation, and the distinction of objects of similar look. In Figure ROCKET-2: Steering Visuomotor Policy via Cross-View Goal Alignment, we show an agent in the left corner and a human player standing on a farmland. The human intends to command the agent to hunt a sheep near the house, even though the agent cannot initially observe the target sheep. To achieve this, the agent must establish spatial relationships using shared visual landmarks between the human’s camera view and its own. We find that relying solely on a behavior cloning loss is insufficient.
@@
To address these challenges, we highlight an important property of behavior datasets (Cai et al., 2025a): The target object remains consistent across camera views in an interaction trajectory. Motivated by this, we propose two auxiliary objective functions: cross-view consistency loss and target visibility loss, to explicitly enhance the agent’s ability to align goals across camera views. Specifically, cross-view consistency loss requires the agent to accurately predict the target object’s centroid point w.r.t. its camera view, while target visibility loss helps the agent determine whether the target object is occluded. By combining these auxiliary losses with behavior cloning loss, we develop ROCKET-2, a state-of-the-art agent in Minecraft. Our experiments show that ROCKET-2 can autonomously track the target object as the camera changes, eliminating the need for SAM’s (Ravi et al., 2024) real-time semantic segmentation, speeding up inference 3×3$\times$3 × to 6×6$\times$6 × compared to ROCKET-1. We observe that ROCKET-2 can interpret intentions from a human’s camera view and make decisions to achieve expected goals in the 3D world. Notably, by simply mapping Minecraft’s action space to that of other 3D environments, such as Doom (Kempka et al., 2016), DMLab (Beattie et al., 2016), and Unreal Engine (Zhong et al., 2024), we are surprised to find that ROCKET-2 could successfully perform basic navigation and object interaction tasks in a zero-shot generalization manner. This demonstrates that our proposed cross-view alignment scheme is domain-agnostic and holds large potential for further scaling across a wider range of environments.
@@
Our contributions are threefold: (1) We introduce a user-friendly interface that allows humans to specify goals using instance-segmentation from humans’ camera view. (2) By introducing cross-view consistency loss and target visibility loss, we train ROCKET-2, an agent that autonomously tracks targets, eliminating the need for real-time goal segmentation and significantly speeding up inference. (3) We demonstrate that ROCKET-2, trained solely on Minecraft data, can zero-shot generalize to other 3D environments, which opens new avenues for discovering universal decision representations in 3D environments.
@@ 2 Related Works
We address policy learning in partially observable 3D environments (Savva et al., 2019; Cai et al., 2024), where the agent perceives only egocentric views and must actively explore to locate key objects (Pearce and Zhu, 2022). Some methods leverage 3D point clouds for global context (Huang et al., 2023; Jiang et al., 2024), but such data is often unavailable. More commonly, memory-based architectures are used to aggregate past observations: RNNs help avoid redundant exploration (Zhao et al., 2023; Gadre et al., 2022), and TransformerXL enables long-horizon planning in Minecraft (Baker et al., 2022; Guss et al., 2019). To mitigate partial observability, some works (Krantz et al., 2023, 2022) use instance images to specify goals, but these require unoccluded, centered views where the object occupies most of the image—constraints that limit their applicability. Instead, we require the policy to align object references across egocentric views, enabling robust target tracking under camera motion. While cross-view alignment has been explored in computer vision for BEV segmentation (Borse et al., 2023) and re-identification (Xu et al., 2019), we are the first to apply it to open-world policy learning.
@@
GCIL optimizes conditional policies via behavior cloning (Pomerleau, 1988), using goal representations such as language (Brohan et al., 2022, 2023; Lynch et al., 2023), images (Majumdar et al., 2022; Lifshitz et al., 2023; Sundaresan et al., ), videos (Cai et al., 2023, 2025b), or trajectory sketches (Wang et al., 2023a; Gu et al., 2023). Compared to standard imitation learning, GCIL provides clearer targets, simplifies behavior modeling, and improves policy controllability. Language goals are commonly used (Padalkar et al., 2023; Driess et al., 2023; Wang et al., 2023d) but often fail to convey spatial details (Cai et al., 2025a; Gu et al., 2023). Image goals (Majumdar et al., 2022; Wang et al., 2023a) better ground spatial intent but are sensitive to lighting, textures, and other irrelevant features. Sketches (Sundaresan et al., ) reduce visual sensitivity but are difficult to generate. Trajectory sketches (Gu et al., 2023) improve control and generalization, but assume full observability, limiting their use in 3D worlds. We instead propose aligning goal-relevant objects across views, enabling spatially grounded yet robust goal conditioning in partially observable environments.
Figure 2: Overview of our approach. We address the challenge of steering visuomotor policy via cross-view goal alignment. Our method allows humans to specify goals from their camera view while the agent acts based on its own observations.
3 Method
@@
Our goal is to learn a goal-conditioned visuomotor policy, which allows humans to specify goal objects for interaction using semantic segmentation across camera views. Formally, we aim to learn a policy $\pi_{\text{cross}}(a_{t}|o_{1:t},\{o_{g},m_{g}\},c_{g})$, where $a_{t}$ represents the action at time $t$, $c_{g}$ denotes the interaction event, such as break item, kill entity, and approach. In the Minecraft environment, an action corresponds to raw mouse and keyboard inputs. $o_{t}\in\mathbb{R}^{H\times W\times 3}$ denotes the environment observation at time $t$, and $o_{g}\in\mathbb{R}^{H\times W\times 3}$ represents an observation of the local environment from a specific camera view. Generally, $o_{g}$ and $o_{t}$ have some visual content overlap. $m_{g}\in\{0,1\}^{H\times W\times 1}$ is a segmentation mask for $o_{g}$, highlighting the target object within the camera view $o_{g}$. During inference, users select a view $o_{g}$ containing the desired object from historical observations returned by the environment and generates its corresponding semantic segmentation $m_{g}$. To train such visuomotor policy, we assume access to a dataset $\mathcal{D}_{\text{cross}}=\{c^{n},(o_{t}^{n},a_{t}^{n},m_{t}^{n})_{t=1}^{L(n)}\}_{n=1}^{N}$ consisting of $N$ successful demonstration episodes, $L(n)$ is the length of episode $n$. Within each episode, if $m_{t}$ is non-empty, all $(o_{t},m_{t})$ pairs indicate the same object. Consequently, we can arbitrarily pick one observation frame as the goal view condition for the entire trajectory.
@@
Without loss of generality, we use the Minecraft world as an example to illustrate the data generation process. Manually collecting datasets that meet the requirements is highly expensive. Thus, we employ the backward trajectory relabeling technique proposed in Cai et al. (2025a) to automate the annotation of the OpenAI Contractor Dataset (Baker et al., 2022), which consists of free-play trajectories from human players: $\mathcal{D}_{\text{raw}}=\{(o_{t}^{n},a_{t}^{n})_{t=1}^{L(n)}\}_{n=1}^{N}$. Specifically, for any given episode $n$, we first detect all frames $o^{n}_{j}$ where interaction events occur, identify the interaction type $c^{n}_{j}$, and localize the interacted object near frame $j$ using bounding boxes and point-based prompts. The SAM-2 (Ravi et al., 2024) model is then employed to generate the segmentation mask $m^{n}_{j}$ for the object. Starting from frame $j$, we traverse the trajectory backward and use the SAM-2 model to continuously generate segmentation masks for the object in real-time until either a new interaction event is encountered or a maximum tracking length is reached. Let $i$ denote the end frame. The resulting trajectory clip is then added to the training dataset: $\mathcal{D}_{\text{cross}}\leftarrow\mathcal{D}_{\text{cross}}\cup\{c_{j},(o^{n}_{t},a^{n}_{t},m^{n}_{t})_{t=i}^{j}\}.$ This ensures that every extracted clip is associated with a consistent interaction intent. The generated data encompasses the fundamental interaction types in Minecraft, including use, break, approach, craft, and kill entity. Among these, approach is a unique event, identified by detecting trajectory clips where the displacement exceeds a specified threshold. The object located at the center of the clip’s final frame is designated as the goal of the approach event.
@@
It is worth noting that incorporating the historical observations $o_{1:t-1}$ as input is essential, especially when there is limited shared visual content between $o_{t}$ and $o_{g}$. This historical sequence acts as a smooth bridge to facilitate alignment. Since the goal object represented by the mask corresponds to the target of the policy’s interaction, this auxiliary task aligns the policy’s actions with its visual focus, effectively improving task performance.
@@
Due to the partial observability in 3D environments, it is common for target objects in interaction trajectories to disappear from the field of view and reappear later. During such intervals, the segmentation mask for the missing object is empty. To leverage this information, we propose training the model to predict whether the target object is currently visible, formulated as: $\pi_{\text{cross}}(v_{t}\mid o_{1:t},\{o_{g},m_{g}\})$, where $v_{t}$ is a binary indicator for empty segmentation masks. On the one hand, accurately predicting object visibility helps the policy better match the target object, avoiding a simple appearance similarity measurement between two frames. On the other hand, visibility information guides the policy to make reasonable decisions, such as confidently approaching the goal when it is visible or actively adjusting its camera to explore when the target is absent.
@@
Let a training trajectory $n$ be denoted as $(c_{g},\{o_{t},m_{t}\}_{t=1}^{L(n)})$. A cross view index $g$ is sampled from $\{i|i\in[1,L(n)],m_{i}\neq\phi\}$. We resize all visual observations and their segmentation masks to $224\times 224$. For encoding the visual observation $o_{t}$, we utilize a DINO-pretrained (Caron et al., 2021) 3-channel ViT-B/16 (Dosovitskiy et al., 2020) (16 is the patch size), which outputs a token sequence of length 196, denoted as $\{\hat{o}_{t}^{i}\}_{i=1}^{196}$. Inspired by Zhong et al. (2025), we encode the segmentation mask $m_{t}$ using a 1-input-channel ViT-tiny/16, yielding $\{\hat{m}_{t}^{i}\}_{i=1}^{196}$. The ViT-base/16 encoder is frozen during training for efficiency, while the ViT-tiny/16 is trainable. To ensure spatial alignment, we fuse the cross-view condition $(o_{g},m_{g})$ by concatenating the feature channels:
@@
Given the ability of self-attention mechanisms to capture spatial details across views, we concatenate the token sequences from two views into a sequence of length 392. A non-causal Transformer encoder module is applied (Vaswani et al., 2017) for spatial fusion, obtaining a frame-level representation $x_{t}$:
@@
Then, we leverage a causal TransformerXL (Dai et al., 2019) architecture to capture temporal information among frames:
@@
Finally, a light network maps $f_{t}$ to predict action $\hat{a}_{t}$, centroid $\hat{p}_{t}$, and visibility $\hat{v}_{t}$. The loss function for episode $n$ follows:
@@
Figure 3: The Evaluation Metric is Spatial-Sensitive. $\checkmark$ and $\times$ indicate the correct and incorrect objects for interaction, respectively. None of the task is trained.
Figure 4: Minecraft Interaction Benchmark Results. Our ROCKET-2 achieves 3×3$\times$3 × to 6×6$\times$6 × faster inference, matching or surpassing ROCKET-1’s peak performance in most tasks. Numbers in parentheses indicate Molmo invocation interval (steps); larger values mean higher FPS. + track denotes real-time SAM-2 tracking between Molmo calls.